1. As an oil well is drilled, each new section of drill pipe supports its own weight and the weight of the pipe and the drill bit beneath it. Calculate the stretch in a new 8.00-m-long steel pipe that supports a 200000-g drill bit and a 5.00-km length of pipe with a linear mass density of 20.0 kg/m. Treat the pipe as a solid cylinder with a 8.00-cm diameter.
  2. A large uniform cylindrical steel rod of density ρ = 8.8 g/cm3 is 4.0 m long and has a diameter of 5.0 cm. The rod is fastened to a concrete floor with its long axis vertical. What is the normal stress in the rod at the cross section located at (a) 2.0 m from its lower end? (b) 2.5m from the lower end?
  3. Two children push on opposite sides of a door during play. Both pushes horizontally and perpendicular to the door. One child push with a force of 20.5 N at a distance of 0.600 m from the hinges and the second child pushes at a distance of 0.650 m. What force must the second child exert to keep the door from moving? Assume friction is negligible.
1
Expert's answer
2021-08-11T09:54:31-0400

1)


"m=(8+5000)20+200=100360\ kg\\F=(100360)(9.8)=983528\ N
\\ \Delta L=4\frac{983528}{1.6\cdot10^{10}(\pi (0.08)^2)}8=9.78\ cm"

2)


"p=(8800)(9.8)(4)=3.4\cdot10^5\ Pa"

3)


"F=\frac{0.6}{0.65}20.5=18.9\ N"